OpenStudy (anonymous):

What is the exact value of cos 60° as found on the unit circle?

OpenStudy (alexis55):

the additional angle is 30 degrees and by the triangle sin(30) = x and cos(60) = x sin(30) = cos(60) cos(60) = cos(2*30) = 1 - 2sin^2(30) sin(30) = 1 - 2sin^2(30) 2sin^2(30) + sin(30) - 1 = 0 (2sin(30) - 1)(sin(30) + 1) = 0 sin(30) = 1/2 --> cos(60) = 1/2 I believe this correct

OpenStudy (anonymous):